OpenStudy (anonymous):

What is the sum? –24 + 32 A. –56 B. –8 C. 8 D. 56

OpenStudy (anonymous):

have you tried solving it?

OpenStudy (anonymous):

which one do you think is the answer?

OpenStudy (anonymous):

\[32-24=?\]

OpenStudy (anonymous):

umm B or c

OpenStudy (falling_in_katt):

c

OpenStudy (anonymous):

i think B

OpenStudy (anonymous):

What is 32-24=?

OpenStudy (falling_in_katt):

8

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

you shouldn't always be giving direct answers @Falling_In_Katt

OpenStudy (anonymous):

@shotgun wrong, why would you put a negative sign if 32 is larger than 24?

OpenStudy (falling_in_katt):

Okay...Sorry @proud_yemeniah_

OpenStudy (anonymous):

Simply put this in a calculator: 32-24

OpenStudy (anonymous):

ohh ok can u help me with 4 more

OpenStudy (anonymous):

Post a question.

OpenStudy (anonymous):

What is the value of d + e + f when d = 20, e = –4 and f = –2? A. –26 B. –22 C. 14 D. 18